The pressure of blood in a man’s abdomen 1.2m above ground is 620mmHg, what will be the velocity of the blood in the abdomen if his heart is 2m above the ground and pumps blood at pressure of 750mmHg.Take density of blood as 0.85kgm-3, g=10ms-2
1
Expert's answer
2021-03-04T17:24:22-0500

"\rho g(h_2-h_1)=p_1-p_2+0.5\rho v^2\\
9.8(2-1.2)=-\frac{133.322}{850}(750-620)+0.5 v^2\\v=5.0\frac{m}{s}"